Education researchers estimate that a student’s score y on the PLAANET exam and the number of days x spent preparing for it are related by the linear model y = 3x + 11 + where is normally distributed with a mean of 0 and standard deviation of 2. Find the probability that a student who studies 12 days will make a score less than 50.
Answer #1

here mean value of score for x=12 is 3*12+11=47

hence from normal distribution:

probability that a student who studies 12 days will make a score less than 50 =P(X<50)

=P(Z<(50-47)/2)=P(Z<1.5)=0.9332